OpenStudy (anonymous):

cos(╥ /4)cos(╥ /6)=?

mathslover (mathslover):

\(\pi=180 \) \(\cos\frac{\pi}{4}=cos\frac{180}{4}=cos45\) do you know what is cos 45 ?

mathslover (mathslover):

ok just remember the tables : cos 45 = \(1\over2\) similarly cos\((\frac{\pi}{6})\) = \(cos{\frac{180}{6}}\) = \(cos30 = \frac{\sqrt{3}}{2}\) hence \(\frac{\sqrt{3}}{2}*\frac{1}{2}=\frac{\sqrt{3}}{4}\)